OpenStudy (kuvyojhmoob):

A college student has finished 52 credit hours with a GPA of 2.8. To get into the program she wishes to enter, she must have a GPA of 3.2. How many additional credit hours of 4.0 work will raise her GPA to 3.2?

OpenStudy (anonymous):

My GPA < 2

OpenStudy (mathstudent55):

You need to do an average of all credits in one calculation. You can't do an average of averages.

OpenStudy (mathstudent55):

Let x = number of additional credits the student needs to get 4.0 GPA on. 52 credits with a GPA of 2.8 means a total of 52 * 2.8 = 145.6 \(\dfrac{145.6 + 4x}{52 + x} = 3.2\) Solve the equation for x.

OpenStudy (kuvyojhmoob):

X=26

OpenStudy (mathstudent55):

correct